OpenStudy (anonymous):

Help Ill reward with medals ! (: Solve the following equation, then complete the statement (just below it) regarding the solution. | x + 9 | = 13 x = _____ or x = _____

OpenStudy (larseighner):

|x+9| = 13 means x + 9 = 13 OR x + 9 = -13

OpenStudy (anonymous):

it says to type my answer like this ? it says my answer is wrong (Type the answers separated by a comma, from smallest to greatest like this example: -15, 10.) @LarsEighner

OpenStudy (larseighner):

Whoops! That should be x+9 = 13 or -(x+9) = 13 in the negative case both signs have to change x = 4 OR x = -22 so ( -22, 4)

OpenStudy (anonymous):

-22 and 4.

OpenStudy (anonymous):

thank you!*(: do you know this one? i am having a hard time with it Solve the following equation, then complete the statement (just below it) regarding the solution. |6x| – 8 = 22 x = _____ or x = _____ (Enter your answers separated by a comma, from least to greatest like this example: -1, 9.) @LarsEighner

OpenStudy (anonymous):

x=5 and -5

OpenStudy (anonymous):

-5,5

OpenStudy (anonymous):

@lacrosseplayer thank you!

OpenStudy (anonymous):

@lacrosseplayer Solve the following equation. |2x| – 6 = 4 (Enter your answers separated by a comma, from least to greatest like this example: -15, 10.) Solve the following equation. |-x| + 4 = 7 (Enter your answers separated by a comma, from least to greatest like this example: -15, 10.)

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

okay thank it i just have a really hard time with this for some reason math is not my strong point

OpenStudy (anonymous):

1) -5,5 2) -3,3

OpenStudy (anonymous):

@lacrosseplayer Solve the following inequality, then complete the expression (just below it) regarding the solution set. 2|x – 3| > 8 x < _____ or x > _____

OpenStudy (anonymous):

x>7

OpenStudy (anonymous):

YOUR A LIFE SAVER THANK YOU @lacrosseplayer

OpenStudy (anonymous):

thanks. but i'm not sure about what x< ?

OpenStudy (anonymous):

@Lovee_Stasiaa

OpenStudy (anonymous):

are there any other problems you need help with?

OpenStudy (anonymous):

ya i think so lol @lacrosseplayer

OpenStudy (anonymous):

i'll be around. just mention me when you need me